If the sides of a triangle are in the ratio 3:4:5. what is the length of each side, if the perimeter of the triangle is 66cm?
Answer by checkley77(12569)

3X+4X+5X=66
12X=66
X=66/12
X=5.5 IS THE MULTIPLIER FOR EACH OF THE RATIOS.
3*5.5=16.5
4*5.5=22
5*5.5=27.5
PROOF:
16.5+22+27.5=66
66=66